Dehner Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on direct labor-hours. The company based its predetermined overhead rate for the current year on the following data:

Total direct labor-hours 64,000

Total fixed manufacturing overhead cost $ 358,400

Variable manufacturing overhead per direct labor-hour $ 5.00

Recently, Job P951 was completed with the following characteristics:

Number of units in the job 20

Total direct labor-hours 100

Direct materials $ 730

Direct labor cost $ 6,400

The unit product cost for Job P951 is closest to: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Total variable overhead estimated=(64000*5)=$320,000

Total overhead estimated=Total variable overhead estimated+Total fixed overhead estimated

=320,000+358400

=$678400

Predetermined overhead rate=Total overhead estimated/Estimated direct labor hours

=678400/64000

=$10.6 per direct labor hour

Hence applied overhead=Predetermined overhead rate*Actual direct labor hours

=10.6*100=$1060

Add:Direct materials=$730

Add:Direct labor=$6400

Total cost=$8190

Hence unit product cost=(8190/20)

=$409.5